OpenStudy (anonymous):

If two numbers have arithmetic mean 2700 and harmonic mean 75, find their geometric mean.

OpenStudy (jack1):

does that seem correct @vinhu ?

OpenStudy (anonymous):

Not quite ! Since the answer is less than 999

OpenStudy (jack1):

might call in the big guns, hang on @Jhannybean and @oldrin.bataku can you guys help...?

OpenStudy (yrelhan4):

let the two numbers be a,b.. arithmetic mean=a+b/2.. geometric mean=sqrt(ab).. two equations, two variables.. find a and b first..

OpenStudy (yrelhan4):

once you have the two numbers.. (1/a + 1/b)/2 gives you the harmonic mean...

OpenStudy (jack1):

ah, cheers @yrelhan4

OpenStudy (yrelhan4):

:)

OpenStudy (yrelhan4):

and.. i should also mention it here, arith mean*geom mean is not equal to harmonic mean.. but for n=2. harmonic mean=G^2/A

OpenStudy (jack1):

we have n=2 here, so doesnt that mean that sqrt (harmonic x Arith) = geometric...?

OpenStudy (yrelhan4):

geometric^2 on the rhs.

OpenStudy (jack1):

sqrt on the LHS...

OpenStudy (yrelhan4):

ah, missed it. my bad. yes thats right.

OpenStudy (jack1):

so my initial guess was correct, and we dont have to solve for a and b...?

OpenStudy (yrelhan4):

no you dont, if you know this formula. but again, the best and the safest thing is to start from the basics.. these formulas are certainly not for beginners.. and i thought you wrote, arith mean*geom=harmonic mean or something like that. sorry about that.

OpenStudy (yrelhan4):

and i hope @vinhu got it?

OpenStudy (jhannybean):

O_o

OpenStudy (jack1):

no, good point, how you get it is important but with this: "once you have the two numbers.. (1/a + 1/b)/2 gives you the harmonic mean..." i think harmonic is n/(1/a + 1/b), so it should be 2 / (1/a + 1/b)... yeah?

OpenStudy (yrelhan4):

err wth. i am really messing it up today. yes. 1/H=(1/a + 1/b)/2.. so H=2/(1/a + 1/b).

OpenStudy (jack1):

cool, thanks man, i learnt something new just then, cheers @yrelhan4

OpenStudy (yrelhan4):

I was wrong twice. thanks to you too. :)

OpenStudy (anonymous):

Could anyone summarize the point for me since I haven't learned this kind of mathematics before !

OpenStudy (yrelhan4):

ok look.. arithmetic mean for two numbers= a+b/2 geometric mean= sqrt(ab) harminic mean= 2/(1/a+1/b).. you are given the values of arith and geom mean.. solve those two equations and find the values of a and b.. then put those values in the third equation to get your answer.

OpenStudy (anonymous):

I'm still in the middle of nowhere. More detail please!!! My a , b is extremely huge. Also they are decimal numbers !

OpenStudy (yrelhan4):

Just round them off to up to one decimal points.. and you have all the details you need to solve this.

OpenStudy (jack1):

a= 37.7 ish, b = 5362.2 ish... yeah? @vinhu

OpenStudy (jack1):

all u need so do is solve them like a series of simultaneous equations arithmetic mean for two numbers= a+b/2 2700 = a + b / 2 5400 = a + b so a = 5400 - b plug that into: harmonic mean= 2/(1/a+1/b) 75 = 2/(1/a+1/b) and a = 5400 - b so 75 = 2/(1/(5400 - b)+1/b) solve for b use that to find a use a and b to find geometirc mean ( sqrt (axb))

hartnn (hartnn):

just an alternate way....might be easier. 5400 = a+b 75= 2ab/ (a+b) = 2ab/ 5400 so, ab = 75*5400/2 = 202500 and Geometric mean is just \(\sqrt{ab}=\sqrt{202500}=450\) see, not much solving was involved :)

hartnn (hartnn):

and definitely no need to find a and b.